# ©Plug-in Authorization for Human Content Copyright Protection in Text-to-Image Model

### v1

@@ Abstract
This paper addresses the contentious issue of copyright infringement in images generated by text-to-image models, sparking debates among AI developers, content creators, and legal entities. State-of-the-art models create high-quality content without crediting original creators, causing concern in the artistic community. To mitigate this, we propose the ©Plug-in Authorization framework, introducing three operations: addition, extraction, and combination. Addition involves training a ©plug-in for specific copyright, facilitating proper credit attribution. Extraction allows creators to reclaim copyright from infringing models, and combination enables users to merge different ©plug-ins. These operations act as permits, incentivizing fair use and providing flexibility in authorization. We present innovative approaches, “Reverse LoRA” for extraction and “EasyMerge” for seamless combination. Experiments in artist-style replication and cartoon IP recreation demonstrate ©plug-ins’ effectiveness, offering a valuable solution for human copyright protection in the age of generative AIs.
keywords:
Copyright Authorization, Copyright Protection, Text-to-image Model, Generative AI
1 Main
Figure 1: (a): ©Plug-in Authorization Process. The authorization process consists of three types of entities: user, model provider, and IP owners (artists). Users can generate copyrighted images only by accessing the relevant plug-in. The model provider offers services to users, tracks usage of plug-ins, and attributes rewards to the IP owner. The IP owners can achieve authorization by registering their ©plug-ins through addition or extraction. These ©plug-ins form a pool where users can use ©plug-ins to produce content with the IP owner’s authorization. (b): Three foundational operations achieving ©Plug-in Authorization: addition, extraction, and combination. The plug-in can be created by addition if the copyrighted work is new to the base model. Meanwhile, the plug-in can be created by extracting from the base model if the copyrighted work is already infringed by the base model. Once a pool of ©plug-ins is constructed, the combination operation can merge multiple ©plug-ins featuring the generation of multiple concepts and leave a non-infringing model excluding all multiple concepts.
@@ 1 Main
Large foundation models are trained with extensive, high-quality datasets like LAION [1] and other enormous undisclosed data sources, which definitely contain copyrighted human contents. At the same time, these models not only excel at generating content based on user prompts [2, 3, 4], but also have the potential of memorizing the exact training data thanks to the huge capacity in their gigantic numbers of parameters [5].
@@
These have sparked copyright infringement concerns among content providers, artists, and users. A notable instance is the lawsuit filed by The New York Times against OpenAI and Microsoft [6], alleging the unauthorized use of a vast number of articles for the purpose of training automated chatbots. The lawsuit seeks the destruction of the allegedly infringing chatbots and their associated training data. Similar concerns and legal actions are also emerging in the field of text-to-image generation [7].
Although it is still debatable whether using copyrighted material to train machine learning models is allowable due to the fair use exception in copyright laws [6, 7], it’s definitely that these powerful models have disrupted existing reward systems for creative artists and content providers [8]. This paper advocates revisiting the goal of copyright laws [9], which is to protect the original expression of ideas in creative works and promote scientific progress by granting authors exclusive property rights for a limited time. However, generative AI models challenge the proper attribution of rewards to copyright holders, impacting artists and domain experts who may hesitate to contribute to knowledge exchange platforms without reasonable rewards. This could lead to a shortage of fresh data for machine learning.
@@
To address these copyright challenges in generative AI models, we introduce the “©Plug-in Authorization” framework for text-to-image generation systems (see Figure  1(a)), to align with existing Intellectual Property (IP) management practices, allowing copyright holders to register their works as plug-ins and receive rewards for their use. End users pay for generating images with copyrighted concepts, and base model providers profit from plug-in registration and usage. This framework facilitates explicit tracking of copyrighted work usage, ensuring a fair reward system.
@@
Technically to enable an effective and efficient copyright authorization, the plug-ins, as permits, should be easily created by addition if copyrighted works are new to the base models, or by extraction if the copyrighted works are already infringed by the base model. Moreover, the plug-ins should be easily combined, which allows copyright holders to merge multiple plug-ins into a new one or enables end users to generate images with multiple copyrighted works. Meanwhile, for efficient execution, these operations should be implemented as light adaptations to the base model, e.g., parameter-efficient tuning methods or prompt designs. We introduce three foundational operations - addition, extraction, and combination - implemented using the Low-Rank Adaptor (LoRA) method [10]. These operations are essential for realizing the Copyright Plug-in Authorization (Figure 1(b) for an overview).
In summary, we make a conceptual contribution: advocating solving the problem of copyright infringement in foundation models with a ©Plug-in Authorization framework. It can offer a fair and practical solution for the attribution challenge in text-to-image generative models. We further introduce three operations addition, extraction and combination to instantiate the framework with efficient human content copyright authorizations. Technically, we propose a novel “Reverse LoRA” algorithm for extraction. It can effectively extract copyrighted concepts from the base model, achieving competitive performance for concept extraction with flexible plug-ins, which are verified with deleting various artist styles and cartoon IPs. We further propose a novel “EasyMerge” approach for combination. It is a data-free layer-wise distillation approach, which can effectively and efficiently address the challenge of combining multiple LoRA components. These methodological innovations are extensively verified with practical text-to-image generation scenarios.
2 Methods: The ©Plug-in Authorization Framework for Text-to-Image Models
@@
As detailed in the Section 1, we implement the “©Plug-in Authorization” by utilizing the publicly available pretrained diffusion generative model, Stable Diffusion [11], along with LoRA components [10]. It is important to note that our framework is not confined to specific model structures, thereby also compatible with other foundation models such as the GPT series [12]. Additionally, it is capable of synergizing with various light fine-tuning or prompt tuning techniques [13, 14, 15, 16]. In the subsequent sections, we revisit the fundamentals of diffusion generative models and introduce the three basic operations of the framework, along with our innovative algorithms.
@@ 2.1 Preliminary on Diffusion Generative Model
Diffusion models [17, 18, 19] are probabilistic models designed to learn a data distribution. In the forward pass, Gaussian noises are successively added $T$ times to an image $X_{0}$, thereby creating a sequence ${X_{0},...,X_{T}}$ that simulates a Markov process. Conversely, the reverse process trains the model to denoise, effectively emulating the reversal of the Markov Chain. New images are generated by initially sampling random Gaussian noises and then denoising them using the model. Importantly, this process can be conditioned on inputs, such as a prompt text $c$. The denoising process, denoted as ${\Phi}_{(w)}(X_{t},c,t)$, is trained to predict the noise under the textual prompt $c$ at any timestep $t\in[0,T]$, as outlined in (1),
$\mathop{\arg\min}\limits_{w}\ \mathbb{E}_{\epsilon,X,c,t}\|{\Phi}_{(w)}(X_{t},c,t)-\epsilon\|^{2},$ (1)
where $w$ denotes the model parameters that are usually of the form U-Net [11] or DiT [20]. Recent advancements have introduced latent diffusion models as a solution to mitigate the drawbacks associated with evaluating and optimizing models in pixel space, such as low inference speed and high training costs. These latent diffusion models operate within a compressed latent space, exemplified by publicly available models such as the Stable Diffusion Model (SDM), as detailed in Rombach et al. [11]. The SDM architecture consists of a variational autoencoder (VAE) that maps images to latent space, a U-Net that learns the diffusion process, and a CLIP encoder for text embedding. Our work primarily focuses on the attention structure within the U-Net, which has been identified as the most influential component in diffusion models.
To implement the “©Plug-in Authorization”, we incorporate three foundational operations into the Stable Diffusion Model (SDM): addition, which allows copyright owners to add a plug-in for their works; extraction, which enables owners to extract a plug-in from an infringing base model; and combination, which permits users to merge plug-ins for multiple copyrighted concepts. The addition operation employs LoRA components that are added to SDM’s attention matrices, and these are then trained with copyrighted data. While the specifics of the addition operation are covered on existing model-sharing platforms like Civitai [21], our discussion will primarily focus on the extraction and combination operations in the subsequent sections.
@@ 2.2 Extraction: Reverse LoRA
To achieve extraction, we introduce a method known as “Reverse LoRA”. This approach consists of two crucial steps to effectively extract the target copyrighted concept while maintaining the ability to generate other irrelevant concepts. The initial step, termed the de-concept step, involves removing the target concept from the base model. The subsequent step is a counterbalance to the de-concept step, involving relearning the surrounding semantic context, referred to as the re-context step. Figure 2 illustrates these two steps in the context of extracting the target concept “Picasso”.
The de-concept step. Our goal is to extract Picasso-related information from the base model to a copyright plug-in using LoRA ($w_{L}$). This involves identifying the information that represents “Picasso” via an alignment process and documenting the changes in model parameters that occur during the alignment process.
@@
$\mathbb{E}[{\Phi}_{({w})}(\epsilon,c^{*},t)]=\mathbb{E}[{\Phi}_{({w+w_{L}})}(\epsilon,c,t)],$ (2)
where $\Phi$ is the denoising function, $w$ denotes the original network parameter, $w_{L}$ is the LoRA component, $c$ is the prompt “the painting of a building”, $c^{*}$ is the prompt “the painting of a building by Picasso”, $\epsilon$ is the initial noise and $t$ is the sampling timestep.
@@
where $X^{*}$ is the copyrighted image (or the image generated by the infringing model with the prompt “the painting of building by Picasso”), $X^{*}_{t}=\sqrt{\alpha_{t}}X^{*}+\sqrt{1-\alpha_{t}}\epsilon$ is the noisy version of $X^{*}$, $c$ is the prompt without copyright concepts, i.e., “the painting of building”, $w$ is the original network and $w_{L}$ is the LoRA weight. This optimization aims to adjust $w_{L}$ so that it effectively captures the desired information related to the copyright concept.
After adding such a LoRA component, the base model can generate Picasso-style images even when the prompts do not explicitly mention “Picasso”. Hence, the LoRA represents the Picasso style, and $w-w_{L}$ would give us a non-infringing model, which can be thought of as an analogy of a negative LoRA. However, directly using $w-w_{L}$ as the non-infringing model compromises its ability to generate images with surrounding contexts, e.g., “the painting of a building”, as shown in Figure 5 in Appendix B. This observation motivates us to further tune the LoRA with pairs of images and texts of surrounding contexts.
The re-context step. To mitigate the performance degradation of the non-infringing model when generating images with contextually related prompts, we introduce a re-context step after the de-concept step. This step involves fine-tuning the LoRA component with images and textual prompts of surrounding contexts, e.g., “the painting of a building”. To curate the dataset, we randomly generate images with the base model using the contextual prompt “the painting of a building”, while leveraging the negative prompt [22] “Picasso” to steer the generation as far away from the target concept “Picasso” as possible. Specifically, we further optimize ${w}_{L}$ with the objective,

### v2

# ©Plug-in Authorization for Human Copyright Protection in Text-to-Image Model

@@ Abstract
This paper addresses the contentious issue of copyright infringement in images generated by text-to-image models, sparking debates among AI developers, content creators, and legal entities. State-of-the-art models create high-quality content without crediting original creators, causing concern in the artistic community and model providers. To mitigate this, we propose the ©Plug-in Authorization framework, introducing three operations: addition, extraction, and combination. Addition involves training a ©plug-in for specific copyright, facilitating proper credit attribution. The extraction allows creators to reclaim copyright from infringing models, and the combination enables users to merge different ©plug-ins. These operations act as permits, incentivizing fair use and providing flexibility in authorization. We present innovative approaches, “Reverse LoRA” for extraction and “EasyMerge” for seamless combination. Experiments in artist-style replication and cartoon IP recreation demonstrate ©plug-ins’ effectiveness, offering a valuable solution for human copyright protection in the age of generative AIs. The code is available at https://github.com/zc1023/-Plug-in-Authorization.git
1 Introduction
@@
Large foundation models Brown et al. ; Touvron et al. (2023); Radford et al. (2021); OpenAI (2023); Rombach et al. (2022b) are trained with extensive, high-quality datasets like The Pile Gao et al. (2020), C4 Raffel et al. (2020), LAION Schuhmann et al. (2022) and other enormous undisclosed data sources, which definitely contain copyrighted human contents. At the same time, these models not only excel at generating content based on user prompts cha ; OpenAI (2023); Rombach et al. (2022b); Ramesh et al. (2021; 2022), but also have the potential of memorizing the exact training data thanks to the huge capacity in their gigantic numbers of parameters Carlini et al. (2021; 2023a).
@@
Such training procedure and utilization of AI models have sparked copyright infringement concerns among content providers, artists, and users. A notable instance is the lawsuit filed by The New York Times against OpenAI and Microsoft NYT , alleging the unauthorized use of a vast number of articles for the purpose of training automated chatbots. The lawsuit seeks the destruction of the allegedly infringing chatbots and their associated training data. Similar concerns and legal actions are also emerging in the field of text-to-image generation law .
Indeed, these concerns are well justified as these powerful models could disrupt the existing reward system in creative arts, adding anxiety to the content providers and artist community. The proficiency of AI in generating artworks that rival human creations is noteworthy, particularly in its ability to replicate characters from major intellectual properties (IPs). For instance, the use of stable diffusion models Rombach et al. (2022a), combined with controlled generation techniques like ControlNet Zhang et al. (2023b), enables users to effortlessly create well-known characters, such as those from Disney. This ease of replication significantly lowers the barriers to potential copyright infringement, raising concerns about increased piracy risks.
One debating point is whether using copyrighted material to train machine learning models is prohibited by copyright laws. It is known that copyright does not ban all forms of copying or replication due to the fair use doctrine, which allows certain copying and distribution if it can be justified as fair use. It is not clear whether AI companies can successfully argue that their training procedures fall under this ’fair use’ exception in copyright laws NYT ; law . Additionally, academic research is underway to develop methods ensuring AI models do not generate copyrighted concepts, as seen in works like Vyas et al. (2023).
In this paper, we step back and advocate to rethink the motivation for enforcing copyright laws. Copyright is a type of intellectual property that intends to protect the original expression of ideas in creative works, which can include literary, artistic, or musical forms cop . The foundational goal of copyright laws, as stated in Article I, Section 8, Clause 8 of the U.S. Constitution USC , is “To promote the Progress of Science and useful Arts, by securing for limited Times to Authors and Inventors the exclusive Right to their respective Writings and Discoveries”. The primary objective of copyright is to incentivize authors to create new works and to facilitate the dissemination of these works to the public by granting them property rights. However, existing generative AI models present challenges in appropriately attributing proper rewards to the copyright holders, which can significantly impact society. Artists, who depend on attribution for recognition and income, may be affected. Additionally, domain experts contributing to knowledge exchange websites like StackOverflow and Quora might hesitate to provide answers if they do not receive reasonable rewards. This situation could backfire on machine learning, as generative models might soon face a shortage of fresh data due to reduced contributions from these sources.
Figure 1: ©Plug-in Authorization Process. The authorization process consists of three types of entities: user, model provider, and IP owners (artists). Users can generate copyrighted images only by accessing the relevant plug-in. The model provider offers services to users, tracks usage of plug-ins, and attributes rewards to the IP owner. The IP owners can achieve authorization by registering their ©plug-ins through addition or extraction. These ©plug-ins form a pool where users can get ©plug-ins to produce content with the IP owner’s authorization.
Figure 2: Three foundational operations achieving ©Plug-in Authorization: addition, extraction, and combination. The plug-in can be created by addition if the copyrighted work is new to the base model. Meanwhile, the plug-in can be created by extracting from the base model if the copyrighted work is already infringed by the base model. Once a pool of ©plug-ins is constructed, the combination operation can merge multiple ©plug-ins featuring the generation of multiple concepts and leave a non-infringing model excluding all multiple concepts.
To address the attribution challenge in generative AI models, the concept of Stable Attribution Troynikov (2023) has been proposed, aiming to credit artists and share revenue with creators based on their contributions. Specifically, Stable Attribution attempts to trace back an AI-generated image to the most similar examples in the training dataset. However, achieving this with reasonable cost and ensuring fairness is challenging, given the vast size and heterogeneous nature of the training set. Content providers and model owners may have completely different views on the evaluation of the content NYT . While copyright data is unique, its impact on model performance may not be able to be fairly measured, especially when compared to the benefits derived from the vast size of open-domain public data. The value of artistic works might be underestimated by model owners. Addressing these issues is essential to create a balanced ecosystem in generative AI.
@@
In addressing the challenge of practical attribution in generative models, we introduce the “Copyright Plug-in Authorization” framework (see Figure 1), designed to align with existing Intellectual Property (IP) management practices. This framework involves base model providers, like Stability AI, functioning as repositories for copyright plug-ins. Copyright holders, such as artists, can register their works as plug-ins, receiving rewards for their use. End users, in turn, pay for the generation of images involving copyrighted concepts using these plug-ins. This system offers positive incentives for all involved: copyright holders are compensated for their creative contributions, end users can use copyrighted plug-ins without risking infringement, and base model providers profit from plug-in registration and model usage. The framework also facilitates explicit tracking of copyrighted work usage, ensuring a fair and straightforward reward system. By successfully implementing this authorization process, we can enable a more equitable distribution of copyright benefits across the generative model landscape.
@@
Technically to enable an effective and efficient copyright authorization, the plug-ins, as permits, should be easily created by addition if copyrighted works are new to the base models, or by extraction if the copyrighted works are already infringed by the base model. Moreover, the plug-ins should be easily combined, which allows copyright holders to merge multiple plug-ins into a new one or enables end users to generate images with multiple copyrighted works. Meanwhile, for efficient execution, these operations should be implemented as light adaptations to the base model, e.g., parameter-efficient tuning methods or prompt designs.
In this paper, we introduce three foundational operations - addition, extraction, and combination - implemented using the Low-Rank Adaptor (LoRA) method Hu et al. (2022). These operations are essential for realizing the Copyright Plug-in Authorization (Figure 2 for an overview).
It is noteworthy that Civitai (civ, ) represents a commendable attempt to instantiate the addition operation, as users can train and share LoRA components to generate corresponding figures. However, the operations extraction and combination are currently not publicly available and pose greater challenges.
The extraction operation involves separating the generative model into a non-infringing base model and some copyrighted plug-ins. A conventional approach might involve retraining the model from scratch using only non-infringing data, and then applying LoRA with copyrighted data. However, this method is impractical, if not impossible, due to high training costs and complex data-cleaning processes. Alternatively, this paper introduces a “Reverse LoRA” approach to extract a plug-in from an infringing base model. This process begins by capturing the target concept: we LoRA-tune the model on the target concept and then take the negative of the LoRA weights to achieve concept destruction. Then we fine-tune the LoRA on surrounding contexts to repair the non-infringing model’s contextual generation ability. Finally, we reverse the LoRA to be the ©plug-in.
The combination operation entails merging multiple copyrighted plug-ins into a unified one. Simply adding these plug-ins together could lead to unpredictable results due to the correlation among copyrighted plug-ins. In this paper, we have successfully developed a method to fuse multiple components. We introduce “EasyMerge”, a method termed “data-free layer-wise distillation” for the combination process. Inspired by conditional generation in generative models, we utilize a LoRA component designed to learn the layer-wise outputs of ©plug-ins under corresponding conditions. Consequently, the LoRA component can mimic the behavior of these ©plug-ins when subjected to the corresponding conditions, effectively achieving the combination of ©plug-ins.
Our contributions are summarized as follows:
  • Conceptual contribution: A ©Plug-in Authorization framework. We advocate to solve the problem of copyright infringement in foundation models with a ©Plug-in Authorization framework. It can offer a fair and practical solution for the attribution challenge in text-to-image generative models. We further introduce three operations addition, extraction and combination to instantiate the framework with efficient human content copyright authorizations.
  • Technical contribution: A novel “Reverse LoRA” algorithm for extraction. It can effectively extract copyrighted concepts from the base model, achieving competitive performance for concept extraction with flexible plug-ins.
  • Technical contribution: A novel “EasyMerge” approach for combination. It is a data-free layer-wise distillation approach, which can effectively and efficiently address the challenge of combining multiple LoRA components.
The structure of the paper is as follows. Section 2 introduces the “©Plug-in Authorization” framework and delves into the three operations addition, extraction and combination. Section 3 presents experiments to validate the effectiveness of the proposed operations. Section 5 concludes the paper, offering a discussion on the limitations of our work.
2 ©Plug-in Authorization with Addition, Extraction and Combination
@@
As detailed in the Introduction, we implement the “©Plug-in Authorization” by utilizing the publicly available pretrained diffusion generative model, Stable Diffusion (Rombach et al., 2022a), along with LoRA components (Hu et al., 2022). It is important to note that our framework is not confined to specific model structures, thereby facilitating compatibility with other foundation models such as the GPT series (Brown et al., ). Additionally, it is capable of synergizing with various light fine-tuning or prompt tuning techniques (Li & Liang, 2021; Lester et al., 2021; Edalati et al., 2022; Hyeon-Woo et al., 2021). In the subsequent sections, we revisit the fundamentals of diffusion generative models and introduce the three basic operations of the framework, along with our innovative algorithms.
@@ 2.1 Preliminary on Diffusion Generative Model
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) are probabilistic models designed to learn a data distribution. In the forward pass, Gaussian noises are successively added $T$ times to an image $X_{0}$, thereby creating a sequence ${X_{0},...,X_{T}}$ that simulates a Markov process. Conversely, the reverse process trains the model to denoise, effectively emulating the reversal of the Markov Chain. New images are generated by initially sampling random Gaussian noises and then denoising them using the model. Importantly, this process can be conditioned on inputs, such as a prompt text $c$. The denoising process, denoted as ${\Phi}_{(w)}(X_{t},c,t)$, is trained to predict the noise under the textual prompt $c$ at any timestep $t\in[0,T]$, as outlined in (1).
$\mathop{\arg\min}\limits_{w}\ \mathbb{E}_{\epsilon,X,c,t}\|{\Phi}_{(w)}(X_{t},c,t)-\epsilon\|^{2}$ (1)
Recent advancements have introduced latent diffusion models as a solution to mitigate the drawbacks associated with evaluating and optimizing models in pixel space, such as low inference speed and high training costs. These latent diffusion models operate within a compressed latent space, exemplified by publicly available models such as the Stable Diffusion Model (SDM), as detailed in Rombach et al. (2022a). The SDM architecture consists of a variational autoencoder (VAE) that maps images to latent space, a U-Net that learns the diffusion process, and a CLIP encoder for text embedding. Our work primarily focuses on the attention structure within the U-Net, which has been identified as the most influential component in diffusion models.
To implement the “©Plug-in Authorization”, we incorporate three foundational operations into the Stable Diffusion Model (SDM): addition, which allows copyright owners to add a plug-in for their works; extraction, which enables owners to extract a plug-in from an infringing base model; and combination, which permits users to merge plug-ins for multiple copyrighted concepts. The addition operation employs LoRA components that are added to SDM’s attention matrices, and these are then trained with copyrighted data. While the specifics of the addition operation are covered on existing model-sharing platforms like civ , our discussion will primarily focus on the extraction and combination operations in the subsequent sections.
@@ 2.2 Extraction: Reverse LoRA
In our endeavor to achieve extraction, we introduce a method known as “Reverse LoRA”. This approach encompasses two crucial steps to effectively extract the target copyrighted concept while maintaining the ability to generate contextual concepts. The initial step, termed the de-concept step, involves removing the target concept from the base model. The subsequent step is a counterbalance to the de-concept step, involving relearning the surrounding semantic context, referred to as the re-context step. Figure 3 illustrates these two steps in the context of extracting the target concept “Picasso”.
2.2.1 Step1: De-concept
Our goal is to extract Picasso-related information from the base model to a copyright plug-in using LoRA ($w_{L}$). This involves identifying the information that represents “Picasso” via an alignment process and documenting the changes in model parameters that occur during the alignment process.
@@ 2.2.1 Step1: De-concept
$\mathbb{E}[{\Phi}_{({w})}(\epsilon,c^{*},t)]=\mathbb{E}[{\Phi}_{({w+w_{L}})}(\epsilon,c,t)]$ (2)
where $\Phi$ is the denoising function, $w$ denotes the original network parameter, $w_{L}$ is the LoRA component, $c$ is the prompt “the painting of a building”,$c^{*}$ is the prompt “the painting of a building by Picasso”, $\epsilon$ is the initial noise, and $t$ is the sampling timestep.
@@
where $X^{*}$ is the copyrighted image (or generated by the infringing model with the prompt “the painting of building by Picasso”), $X^{*}_{t}=\sqrt{\alpha_{t}}X^{*}+\sqrt{1-\alpha_{t}}\epsilon$ is the noisy version of $X^{*}$, $c$ is the prompt of “the painting of building”, $w$ is the original network and $w_{L}$ is the LoRA weight. This optimization aims to adjust $w_{L}$ so that it effectively captures the desired information related to the target concept.
By incorporating such a LoRA component, the base model can generate Picasso-style images even when the prompts do not explicitly mention “Picasso”. Hence, the LoRA represents the copyrighted Picasso style, and $w-w_{L}$ would give us a non-infringing model, which can thought of as an analogy of a ‘negative LoRA’. However, directly using $w-w_{L}$ as the non-infringing model compromises its ability to generate images with surrounding context, e.g., “the painting of a building”, as shown in Figure 7 in Appendix A.1. This observation leads us to further tune the LoRA with pairs of images and texts of surrounding semantic context.
2.2.2 Step2: Re-context
To mitigate the performance degradation of the non-infringing model when generating images with contextually related prompts, we introduce a re-context step following the de-concept step. This step involves fine-tuning the LoRA component with images and textual prompts of surrounding contexts, e.g., “the painting of a building”. To curate the dataset, we randomly generate images with the base model using the contextual prompt “the painting of a building”, while leveraging the negative prompt (Ho & Salimans, 2022) “Picasso” to steer the generation as far away from the target concept “Picasso” as possible.
Specifically, we further optimize ${w}_{L}$ with the objective,